# On-device Sora: Enabling Diffusion-Based Text-to-Video Generation for Mobile Devices

### v1

@@ Abstract.
We present On-device Sora, a first pioneering solution for diffusion-based on-device text-to-video generation that operates efficiently on smartphone-grade devices. Building on Open-Sora, On-device Sora applies three novel techniques to address the challenges of diffusion-based text-to-video generation on computation- and memory-limited mobile devices. First, Linear Proportional Leap (LPL) reduces the excessive denoising steps required in video diffusion through an efficient leap-based approach. Second, Temporal Dimension Token Merging (TDTM) minimizes intensive token-processing computation in attention layers by merging consecutive tokens along the temporal dimension. Third, Concurrent Inference with Dynamic Loading (CI-DL) dynamically partitions large models into smaller blocks and loads them into memory for concurrent model inference, effectively addressing the challenges of limited device memory. We implement On-device Sora on the iPhone 15 Pro, and the experimental evaluations demonstrate that it is capable of generating high-quality videos on the device, comparable to those produced by Open-Sora running on high-end GPUs. These results show that On-device Sora enables efficient and high-quality video generation on resource-constrained mobile devices, expanding accessibility, ensuring user privacy, reducing dependence on cloud infrastructure, and lowering associated costs. We envision the proposed On-device Sora as a significant first step toward democratizing state-of-the-art generative technologies, enabling video generation capabilities on commodity mobile and embedded devices. The code implementation is publicly available at an GitHub repository [^1] https://github.com/eai-lab/On-device-Sora.
copyright: none
1. Introduction
@@
Recent advancements in generative models (Suzuki and Matsuo, 2022) have significantly expanded capabilities in data generation across various modalities, including text (Hu et al., 2017), image (Shaham et al., 2019), and video (Unterthiner et al., 2018). In particular, diffusion-based models for image tasks (Ho et al., 2020; Song et al., 2020; Peebles and Xie, 2023; Rombach et al., 2022; Saharia et al., 2022; Dhariwal and Nichol, 2021; Zhang et al., 2023a; Podell et al., 2023) have emerged as foundational tools for a wide range of applications, such as image generation (Ho et al., 2022b), image editing (Kawar et al., 2023), and personalized content creation (Zhang et al., 2024). Further extending these technologies, diffusion-based generative models are now driving remarkable progress in video generation tasks (Singer et al., 2022; Menapace et al., 2024; Ho et al., 2022a, c; Blattmann et al., 2023a; Molad et al., 2023; Harvey et al., 2022; Blattmann et al., 2023b; Voleti et al., 2022; Wu et al., 2023; Khachatryan et al., 2023), including video synthesis (Liu et al., 2021a) and real-time video analysis (Orriols, 2004), facilitating the development of advanced generative video services and applications in new and unprecedented ways.
One of the most notable video generation models is Sora (Liu et al., 2024), which has demonstrated remarkable potential in creating high-quality videos from textual descriptions (prompts), effectively transforming abstract concepts into visually realistic content. However, the inherently high complexity of diffusion processes for video generation, combined with the substantial size of diffusion-based Transformer models (Peebles and Xie, 2023), imposes significant computational and memory demands. For instance, Sora (Liu et al., 2024) is estimated to take an hour to produce five minutes of video using an NVIDIA H100 GPU of 80 GB RAM. Compared to large language models (LLMs) (Zhao et al., 2023b), the costs for diffusion-based video generation models like Sora are several orders of magnitude higher, indicating their resource-intensive nature. As a result, the accessibility of video generative services remains predominantly confined to environments with extensive computational resources, limiting their applicability in resource-constrained devices.
With the rapid expansion of mobile, IoT, and embedded devices, there is a growing demand for running generative applications directly on the device. In response, numerous studies have primarily focused on the development of on-device image generation (Li et al., 2024b; Vasu et al., 2023; Castells et al., 2024; Choi et al., 2023; Chen et al., 2023; Zhao et al., 2023a), as image generation is comparatively less resource-intensive than video. However, enabling on-device video generation remains in its nascent stages due to the significantly higher computational and memory resource demands compared to image generation (Elasri et al., 2022). Moreover, given that video data is typically much larger and more sensitive than image data (Deng et al., 2023), there is an increasing need for innovative technologies that can enable efficient and effective on-device video generation.
@@
We introduce On-device Sora, the first standalone framework for diffusion-based on-device text-to-video generation, which is capable of producing high-quality videos on smartphone-grade devices. On-device Sora leverages Open-Sora (Zheng et al., 2024) as its backbone and significantly enhances the efficiency of video generation while maintaining the comparable video quality, enabling on-device text-to-video generation with limited computational power and memory. To the best of our knowledge, On-device Sora proposed in this work is the first solution that enables the efficient generation of Sora-level video directly on the device, which has previously been restricted to large-scale data center environments.
@@
On-device Sora addresses three fundamental challenges in enabling diffusion-based on-device text-to-video generation. First, diffusion-based video generation entails a substantial number of denoising steps, requiring repeated executions of diffusion models, ranging from dozens to thousands of iterations. To address this, we propose Linear Proportional Leap (LPL), which reduces nearly half of the denoising steps by leaping through a large portion of steps using the Euler’s method (Biswas et al., 2013) in an estimated direct trajectory to generate videos efficiently. Second, state-of-the-art diffusion models, such as STDiT (Spatial-Temporal Diffusion Transformer) (Zheng et al., 2024), employed by Open-Sora, exhibit high computational complexity due to their substantial amount of token processing (Bolya et al., 2022) in attention modules (Vaswani, 2017). To tackle the high computational complexity of STDiT, we propose Temporal Dimension Token Merging (TDTM), which merges consecutive tokens (Bolya et al., 2022, 2023) in temporal order at attention layers of STDiT, effectively reducing the token processing up to a quarter. Lastly, the substantial size of state-of-the-art models required for text-to-video generation, i.e., T5 (Raffel et al., 2020) and STDiT (Zheng et al., 2024), presents challenges for execution on memory-constrained mobile devices. To overcome this, we propose Conference Inference and Dynamic Loading (CI-DL), which integrates concurrent model inference with the dynamic loading of models into memory. It dynamically divides the model into blocks and loads them into memory based on the device’s available memory capacity at run-time, allowing simultaneous model block loading and inference. With these three proposed methods, high-quality video generation becomes feasible on smartphone-grade devices with limited computing resources, overcoming the requirements for substantial computational power, such as high-end GPUs.
Currently, video generative technology remains neither widely accessible nor commonly available to the public (Liu et al., 2024). Therefore, enabling video generation on commodity mobile and embedded devices can not only enhance the accessibility of advanced video generation technologies but also provide additional benefits. Users can ensure privacy and mitigate concerns about data transmission and leakage of sensitive personal information. On-device video generation eliminates the need for interaction with the third-party cloud servers, which may pose security risks due to potential data breaches or unauthorized access. Furthermore, cloud-based generative models often suffer from latency issues influenced by network speed and may not function optimally when connectivity is unstable or unavailable. On-device generative models, in contrast, can provide more reliable services, offering stable functionality regardless of network conditions. Another limitation of cloud-based video generation is their shared nature, which limits personalization and the ability to cater to individual users. On-device video generation, however, can be tailored to each user, enabling personalized experiences through customized video creations. Through on-device generation, videos can be fine-tuned and customized for individual users, providing contents that better reflect personal preferences, data, contexts, and environments. From a financial perspective, generating videos on-device is far more economical. For example, the cost of an NVIDIA H100 GPU is around \$25,000, whereas an iPhone 15 Pro is about only \$999, which is 25 times more affordable. Additionally, the annual maintenance costs for cloud-based video generation are estimated to range between \$250,000 and \$800,000, whereas a smartphone incurs negligible maintenance expenses.
@@
We implement On-device Sora on the iPhone 15 Pro (Apple, 2023) based on Open-Sora (Zheng et al., 2024), which is an open-source text-to-video generation framework, enabling standalone text-to-video generation directly on the device. The full implementation is available as open-source code in a publicly accessible anonymous GitHub repository1. The extensive experiments are conducted to evaluate the performance of the proposed On-device Sora using the state-of-the-art video benchmark tool, i.e., VBench (Huang et al., 2024), compared with Open-Sora running on NVIDIA A6000 GPUs. The experimental results demonstrate that On-device Sora can effectively generate videos of equivalent quality to Open-Sora (Zheng et al., 2024), while accelerating video generation when applying the proposed three methods. While the iPhone 15 Pro has a GPU of 143 times less computational power (Apple, 2023) and 16 times smaller memory (RAM) compared to the NVIDIA A6000 GPU, the evaluation results demonstrate that On-device Sora significantly improves the efficiency of video generation by effectively compensating for the limited computing resources of the device.
2. Background and Key Challenges
@@
We first provide 1) a background of Open-Sora (Zheng et al., 2024), an open-source video diffusion model, which is the backbone of the proposed On-device Sora, and 2) key challenges in realizing on-device video generation for mobile devices.
2.1. Background: Open-Sora
Figure 1 illustrates the structure of Open-Sora (Zheng et al., 2024), generating videos from prompts (texts) through 1) prompt embedding, 2) latent video generation, and 3) video decoding.
@@ 2.1. Background: Open-Sora
Figure 1. Open-Sora (Zheng et al., 2024) generates realistic videos from the user prompt (text) through three stages: 1) prompt embedding, 2) latent video generation, and 3) video decoding.
@@
1) Prompt (Text) Embedding. The first stage of Open-Sora (Zheng et al., 2024) is to map a user prompt, a textual description of the desired video, to an embedding vector, which is used as input for the subsequent video generation stage. To produce prompt embeddings from user texts, Open-Sora employs T5 (Text-to-Text Transformer) (Raffel et al., 2020), a language model specifically fine-tuned to support video generation tasks.
@@
2) Latent Video Generation. The next stage is to generate the latent video representation conditioned on the prompt embedding obtained from T5 (Text-to-Text Transformer) (Raffel et al., 2020). To this end, Open-Sora employs STDiT (Spatial-Temporal Diffusion Transformer) (Zheng et al., 2024), a diffusion-based text-to-video model using the Markov chain (Zhang et al., 2023b). Since maintaining temporal consistency across video frames is essential in video generation, STDiT (Zheng et al., 2024) applies the spatial-temporal attention mechanism (Yan et al., 2019) to the patch representations. It allows effective learning of temporal features across video frames through the temporal attention, enhanced by incorporating rope embeddings (Su et al., 2024). During the forward process of STDiT, the Gaussian noise $\boldsymbol{\epsilon}_{k}$ is iteratively added to the latent video representation $\mathbf{x}_{k}$ over $K$ steps, transforming the intact video representation $\mathbf{x}_{0}$ into the complete Gaussian noise $\mathbf{x}_{K}$ in the latent space with the forward distribution $q(\mathbf{x}_{k}|\mathbf{x}_{k-1})$ as:
@@
(1) $\begin{split}\mathbf{x}_{k}&=\sqrt{1-\beta_{k}}\mathbf{x}_{k-1}+\sqrt{\beta_{k}}\boldsymbol{\epsilon}_{k}\\ \vspace{-4pt}q(\mathbf{x}_{t}|\mathbf{x}_{k-1})&=\mathcal{N}(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I})\end{split}$
@@
To generate (recover) the latent representation $\mathbf{x}_{0}$, the noise $\boldsymbol{\epsilon}_{k}$ is repeatedly removed (denoised) from the complete Gaussian noise $\mathbf{x}_{K}$ through the reverse process using the estimated noise with the reverse distribution $p_{\boldsymbol{\theta}}(\mathbf{x}_{k-1}|\mathbf{x}_{t})$ modeled by STDiT (Zheng et al., 2024) with the parameter set $\boldsymbol{\theta}$, as follows:
@@
(2) $\begin{split}\mathbf{x}_{k-1}&=\boldsymbol{\mu}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k)+\sqrt{\beta_{k}}\boldsymbol{\epsilon}~{}\text{ where}~{}\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\\ p_{\boldsymbol{\theta}}(\mathbf{x}_{k-1}|\mathbf{x}_{k})&=\mathcal{N}(\mathbf{x}_{k-1};\boldsymbol{\mu}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k),\boldsymbol{\Sigma}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k))\end{split}$
@@
where $\boldsymbol{\mu}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k)$ and $\boldsymbol{\Sigma}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k)$ is the mean and variance of $\mathbf{x}_{k}$ estimated by the STDiT model training, respectively. This reverse process repeatedly denoises $\mathbf{x}_{k}$ into $\mathbf{x}_{k-1}$ over a large number of $1{\leq}k{\leq}K$ denoising steps (i.e., from dozens to thousands of steps), eventually generating the de-noised latent video representation $\mathbf{x}_{0}$ close to the intact representation.
3) Video Decoding. Finally, the latent video representation $\mathbf{x}_{0}$ generated from STDiT is decoded and up-scaled into the human-recognizable video through the VAE (Variational Autoencoder) (Doersch, 2016). The VAE employed in Open-Sora utilizes both 2D and 3D structures; the 2D VAE is based on SDXL (Podell et al., 2023), while the 3D VAE adopts the architecture of Magvit-v2 (Yu et al., 2023).
2.2. Challenges in On-device Video Generation
Although Open-Sora (Zheng et al., 2024) has enabled the generation of high-quality videos from texts, adapting this advanced generative capability to mobile devices presents key challenges.
@@ 2.2. Challenges in On-device Video Generation
Figure 2. On-device Sora enables efficient text-to-video generation directly on the device by employing three key approaches: 1) Linear Proportional Leap, 2) Temporal Dimension Token Merging, and 3) Concurrent Inference with Dynamic Loading.
@@
C1) Excessive Denoising Steps. Table 1 shows the latency of each model component in Open-Sora, where the latent video generation process (denoising process) performed by STDiT is the most time-consuming. That is because a substantial number of denoising steps is required to remove the noise $\boldsymbol{\epsilon}_{k}$ from $\mathbf{x}_{K}$ to obtain $\mathbf{x}_{0}$ during latent video generation (Song et al., 2020). Such extensive denoising iterations presents considerable challenges on mobile devices with constrained computational capabilities. To reduce the number of denoising steps, Open-Sora employs Rectified Flow (Liu et al., 2022), which redefines the objective of diffusion models, inducing the model to directly approximate the straight mapping trajectory between the initial and target distributions. In image generation, it enables a single-step denoising without a significant loss in image quality. However, in video generation models (Abdi, 2007; Zheng et al., 2024), the accumulated variance across sequential video frames prevents the trajectory from being fully represented with a single drift. As a result, Open-Sora (Zheng et al., 2024) requires at least $K=30$ or $50$ denoising steps to produce high-quality videos. While the numerous denoising steps are manageable in server-level environments—where the complete denoising process typically finishes within one minute—on mobile devices, it may take several tens of minutes for full denoising. Accordingly, a new approach to reduce denoising steps is essential to enable efficient on-device video generation.
@@
Table 1. The number of executions (iterations) of each model component (i.e., T5 (Raffel et al., 2020), STDiT (Zheng et al., 2024), and VAE (Doersch, 2016)) in Open-Sora (Zheng et al., 2024) and their total latencies on iPhone 15 Pro (Apple, 2023).
@@
C2) Intensive Token Processing. While a large number of denoising steps poses a significant challenge to video generation on mobile devices, even a single denoising step itself is computationally intensive. The primary reason is that the computational complexity of the attention mechanism (Niu et al., 2021) in STDiT (Zheng et al., 2024) grows quadratically with the token size, which significantly increases the computational load for token processing and, consequently, the model’s inference latency. To address this challenge, Token merging (Bolya et al., 2023) has been proposed to improve the throughput of vision Transformer models. Token merging progressively merges similar visual tokens within the transformer to accelerate the model inference latency by reducing the size of tokens to be processed. While token merging has been applied to diffusion models, it is only applied to spatial tokens (Bolya et al., 2022, 2023) and has not been applied to the temporal tokens in video diffusion models, such as STDiT (Zheng et al., 2024). Thus, a novel token merging method is required to improve the computational efficiency of token processing in video generation, while preserving high video quality.
@@
C3) High Memory Requirements. Figure 3 shows the memory requirements of model components executed by Open-Sora (Zheng et al., 2024), i.e., VAE (Doersch, 2016), T5 (Raffel et al., 2020), and STDiT (Zheng et al., 2024), where their cumulative memory demand, i.e., 23 GB, can easily surpass the memory capacity of many mobile devices. For instance, the iPhone 15 Pro (Apple, 2023), with 8 GB of memory, restricts the available memory for a single application to 3.3 GB to ensure the system stability. Furthermore, the individual memory requirements of T5 and STDiT exceed 3.3 GB, creating challenges in loading them into memory. In addition, some memory must be reserved for model execution (inference), complicating the deployment of Open-Sora on mobile devices. Thus, executing large video generative models with limited device memory is another challenge that should be addressed to enable on-device video generation.
@@
Figure 3. The size of Open-Sora models: T5 (Raffel et al., 2020) (18.00 GB), STDiT (Zheng et al., 2024) (4.50 GB), and VAE (Doersch, 2016) (0.82 GB), which exceeds the available memory capacity of iPhone 15 Pro (Apple, 2023) (3.3 GB).
3. Overview: On-device Sora
Building on Open-Sora (Zheng et al., 2024) outlined above (Sec. 2.1), we propose On-device Sora (Figure 2), which enables diffusion-based text-to-video generation on mobile devices by addressing the key challenges (Sec. 2.2), as summarized below.
1) Linear Proportional Leap (Sec. 4). To reduce the excessive number of denoising steps, we propose Linear Proportional Leap, which enables the generation of high-quality videos with nearly half of the required full denoising steps. This is achieved by leveraging the linear trajectory properties of Rectified Flow (Liu et al., 2022) employed in STDiT (Zheng et al., 2024). Instead of performing the full sequence of denoising steps during video generation, it makes a direct leap along the linear trajectory toward the target data at the middle of the denoising process by utilizing the pre-trained flow fields (Liu et al., 2022). Thus, Linear Proportional Leap effectively reduces the most time-consuming stage of video generation—the iterative execution of STDiT (Zheng et al., 2024)—by minimizing the number of denoising steps. Notably, this reduction in denoising steps is accomplished without necessitating additional model training, architectural modifications, or data calibration, allowing for saving substantial time and computational resources.
2) Temporal Dimension Token Merging (Sec. 5). To lighten the intensive computation required for token processing, we propose Temporal Dimension Token Merging, which merges consecutive video frames in the form of latent representations at each attention layer of STDiT (Zheng et al., 2024). The proposed token merging reduces the amount of tokens to be processed by half and lowers the computational load of attention modules up to one-quarter. To the best of our knowledge, it is the first to merge tokens in temporal order at attention layers in diffusion-based video generation models.
3) Concurrent Inference with Dynamic Loading (Sec. 6). To execute large video generative models (i.e., T5 (Raffel et al., 2020) and STDiT (Zheng et al., 2024)) with the limited device memory, we propose Concurrent Inference with Dynamic Loading, which partitions the models into smaller blocks that can be loaded into the memory and executed concurrently. By parallelizing model execution and block loading, it effectively accelerates iterative model inference, e.g., multiple denoising steps. Also, it improves memory utilization while minimizing the block loading overhead by retaining specific model blocks in memory dynamically based on the available runtime memory.
4. Linear Proportional Leap

### v2

# On-device Sora: Enabling Training-Free Diffusion-based Text-to-Video Generation for Mobile Devices

@@ Abstract
We present On-device Sora, the first model training-free solution for diffusion-based on-device text-to-video generation that operates efficiently on smartphone-grade devices. To address the challenges of diffusion-based text-to-video generation on computation- and memory-limited mobile devices, the proposed On-device Sora applies three novel techniques to pre-trained video generative models. First, Linear Proportional Leap (LPL) reduces the excessive denoising steps required in video diffusion through an efficient leap-based approach. Second, Temporal Dimension Token Merging (TDTM) minimizes intensive token-processing computation in attention layers by merging consecutive tokens along the temporal dimension. Third, Concurrent Inference with Dynamic Loading (CI-DL) dynamically partitions large models into smaller blocks and loads them into memory for concurrent model inference, effectively addressing the challenges of limited device memory. We implement On-device Sora on the iPhone 15 Pro, and the experimental evaluations show that it is capable of generating high-quality videos on the device, comparable to those produced by high-end GPUs. These results show that On-device Sora enables efficient and high-quality video generation on resource-constrained mobile devices. We envision the proposed On-device Sora as a significant first step toward democratizing state-of-the-art generative technologies, enabling video generation on commodity mobile and embedded devices without resource-intensive re-training for model optimization (compression). The code implementation is available at a GitHub repository [^1] https://github.com/eai-lab/On-device-Sora.
1 Introduction
@@
Recent advancements in generative models [60] have significantly expanded capabilities in data generation across various modalities, including text [27], image [55], and video [63]. In particular, diffusion-based models for image tasks [23, 58, 46, 50, 53, 14, 78, 47] have emerged as foundational tools for a wide range of applications, such as image generation [25], image editing [30], and personalized content creation [79]. Further extending these technologies, diffusion-based generative models are now driving remarkable progress in video generation tasks [56, 41, 24, 26, 4, 42, 21, 5, 65, 68, 31], including video synthesis [37] and real-time video analysis [44].
With the rapid expansion of mobile and embedded devices, there is an increasing demand for executing generative applications directly on-device. In response, extensive research has focused on developing on-device image generation methods [35, 64, 8, 13, 12, 81], with recent advancements extending to on-device video generation [71, 70]. However, generating lightweight and compressed on-device video generation models necessitates additional model optimization (training), such as distillation [66, 36, 57], quantization [62, 80], and pruning [69], to reduce model size and computational complexity. Furthermore, mitigating performance degradation in the compressed models requires a time-intensive and iterative model re-training. This process usually demands substantial computational and memory resources, e.g., SnapGen-V [70] takes over 150K training iterations on 256 NVIDIA A100 80GB GPUs to generate mobile-deployable compressed video generation models.
@@
We introduce On-Device Sora, the first training-free framework that enables diffusion-based text-to-video generation on mobile devices, which allows video generative models to run directly on-device, enabling the production of high-quality videos on smartphone-grade devices. While previous approaches [70] require additional training and substantial GPU resources to optimize video generation models for mobile environments, the proposed On-Device Sora eliminates the need for training and directly enables efficient on-device execution. Using pre-trained video generative models, e.g., Open-Sora [83], Pyramidal Flow [29], On-device Sora significantly enhances their efficiency, enabling on-device video generation with limited resources.
@@
To achieve this, we address key challenges of enabling on-device text-to-video generation: 1) Linear Proportional Leap (LPL) reduces the iterative denoising steps in the diffusion process by leaping through a significant portion of steps using the Euler’s method [3] along an estimated direct trajectory, 2) Temporal Dimension Token Merging (TDTM) lowers computational complexity of STDiT (Spatial-Temporal Diffusion Transformer) [83] by merging consecutive tokens [6, 7] in the attention layers, and 3) Conference Inference and Dynamic Loading (CI-DL) enables video generation with limited memory capacity of mobile devices by integrating concurrent model inference with the dynamic loading of models into memory. With these three proposed methods, high-quality video generation becomes feasible on smartphone-grade devices with limited computing resources, overcoming the requirements for substantial computational power, such as high-end GPUs. To the best of our knowledge, On-device Sora proposed in this work is the first training-free solution that enables the efficient generation of video directly on the device. The advantages of On-device Sora, i.e., directly deploying well-trained video generative models onto resource-constrained devices without time-consuming model modification and/or re-training, are expected to become even more pronounced with the active development of compact DiT-based text-to-video generative models [46].
@@
We implement On-device Sora on the iPhone 15 Pro [1] using Open-Sora [83]. The full implementation is available as open-source code in an anonymous GitHub1. The extensive experiments are conducted to evaluate the performance of On-device Sora using the state-of-the-art video benchmark tool, i.e., VBench [28], compared with NVIDIA A6000 GPUs. We also evaluate the proposed methods on additional text-to-video generative models, including Pyramidal Flow [29]. The experimental results demonstrate that On-device Sora maintains comparable video quality while accelerating generation speed with the proposed methods. While the iPhone 15 Pro has a GPU of 143 times less computational power [1] and 16 times smaller memory compared to the NVIDIA A6000 GPU, the evaluation results show that On-device Sora significantly improves the efficiency of video generation by effectively compensating for the limited computing resources of the device.
2 Background and Challenges
@@
We first provide a background of diffusion-based text-to-video generation (e.g. Open-Sora [83]), which is the backbone of the proposed On-device Sora, and key challenges in realizing on-device video generation for mobile devices.
2.1 Background: Diffusion-based Text-to-Video Generation
@@
In general, diffusion-based text-to-video generation models, such as Open-Sora [83], generate videos from user prompts (texts) through the three stages: 1) prompt (text) embedding, 2) latent video generation, and 3) video decoding.
@@
1) Prompt (Text) Embedding. The first stage of diffusion-based text-to-video generation is to map a user prompt, a textual description of the desired video, to an embedding vector, which is used as input for the subsequent video generation stage. For example, to produce prompt embeddings from user texts, Open-Sora employs T5 (Text-to-Text Transformer) [48], a language model specifically fine-tuned to support video generation tasks.
@@
2) Latent Video Generation. The next stage is to generate the latent video representation conditioned on the prompt embedding obtained from language models, e.g., T5 (Text-to-Text Transformer) [48]. To this end, for instance, Open-Sora employs STDiT (Spatial-Temporal Diffusion Transformer) [83], a diffusion-based text-to-video model using the Markov chain [77]. Since maintaining temporal consistency across video frames is essential in video, STDiT [83] applies the spatial-temporal attention mechanism [72] to the patch representations. It allows effective learning of temporal features across video frames through the temporal attention, enhanced by incorporating rope embeddings [59]. During the forward process of STDiT, the Gaussian noise $\boldsymbol{\epsilon}_{k}$ is iteratively added to the latent video representation $\mathbf{x}_{k}$ over $K$ steps, transforming the intact video representation $\mathbf{x}_{0}$ into the complete Gaussian noise $\mathbf{x}_{K}$ in the latent space with the forward distribution $q(\mathbf{x}_{k}|\mathbf{x}_{k-1})$ as:
@@
$\begin{split}\mathbf{x}_{k}&=\sqrt{1-\beta_{k}}\mathbf{x}_{k-1}+\sqrt{\beta_{k}}\boldsymbol{\epsilon}_{k}\\ q(\mathbf{x}_{t}|\mathbf{x}_{k-1})&=\mathcal{N}(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I})\end{split}$ (1)
@@
To generate the latent representation $\mathbf{x}_{0}$, the noise $\boldsymbol{\epsilon}_{k}$ is repeatedly removed (denoised) from the complete Gaussian noise $\mathbf{x}_{K}$ through the reverse process using the estimated noise with the reverse distribution $p_{\boldsymbol{\theta}}(\mathbf{x}_{k-1}|\mathbf{x}_{t})$ modeled by STDiT [83] with the parameter set $\boldsymbol{\theta}$, as follows:
@@
$\begin{split}\mathbf{x}_{k-1}&=\boldsymbol{\mu}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k)+\sqrt{\beta_{k}}\boldsymbol{\epsilon}~{}\text{ where}~{}\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\\ p_{\boldsymbol{\theta}}(\mathbf{x}_{k-1}|\mathbf{x}_{k})&=\mathcal{N}(\mathbf{x}_{k-1};\boldsymbol{\mu}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k),\boldsymbol{\Sigma}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k))\end{split}$ (2)
@@
where $\boldsymbol{\mu}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k)$ and $\boldsymbol{\Sigma}_{\boldsymbol{\theta}}(\mathbf{x}_{k},k)$ is the mean and variance of $\mathbf{x}_{k}$ estimated by STDiT training, respectively. This reverse process repeatedly denoises $\mathbf{x}_{k}$ into $\mathbf{x}_{k-1}$ over a large number of $1{\leq}k{\leq}K$ denoising steps (i.e., from dozens to thousands of steps), eventually generating the de-noised latent video representation $\mathbf{x}_{0}$ close to the intact representation.
@@
Figure 1: On-device Sora enables training-free text-to-video generation directly on the device by employing three key methods: 1) Linear Proportional Leap (LPL), 2) Temporal Dimension Token Merging (TDTM), and 3) Concurrent Inference with Dynamic Loading (CI-DL).
3) Video Decoding. Finally, the latent video representation $\mathbf{x}_{0}$ generated from STDiT is decoded and up-scaled into the human-recognizable video through the VAE (Variational Autoencoder) [15]. For example, the VAE employed in Open-Sora [83] utilizes both 2D and 3D structures; the 2D VAE is based on SDXL [47], while the 3D VAE adopts the architecture of MAGVIT-v2 [76].
2.2 Key Challenges in On-device Video Generation
@@
C1) Excessive Denoising Steps. Tab. 1 shows the latency of each model component in Open-Sora [83], where the latent video generation process (denoising process) performed by STDiT is the most time-consuming. That is because a substantial number of denoising steps is required to remove the noise $\boldsymbol{\epsilon}_{k}$ from $\mathbf{x}_{K}$ to obtain $\mathbf{x}_{0}$ during latent video generation [58]. Such extensive denoising iterations presents considerable challenges on mobile devices with constrained computational capabilities. While the numerous denoising steps are manageable in server-level environments—where the complete denoising process typically finishes within one minute—on mobile devices, it may take several tens of minutes. Accordingly, an effective approach to reduce denoising steps without model modification or re-training is essential to enable on-device video generation.
@@
Table 1: The number of executions (iterations) of each model component (i.e., T5 [48], STDiT [83], and VAE [15]) in Open-Sora [83] and their total latencies on the iPhone 15 Pro [1].
@@
C2) Intensive Token Processing. While a large number of denoising steps poses a significant challenge to video generation on mobile devices, even a single denoising step itself is computationally intensive. The primary reason is that the computational complexity of the attention mechanism [43] in STDiT [83] grows quadratically with the token size, which significantly increases the computational load for token processing and, consequently, the model’s inference latency. To address this challenge, Token merging [7] has been proposed to improve the throughput of vision Transformer models, which progressively merges similar visual tokens within the transformer to accelerate the model inference latency by reducing the size of tokens to be processed. While token merging has been applied to diffusion models, it is only applied to spatial tokens [6, 7] and has not been applied to the temporal tokens in video diffusion models, such as STDiT [83]. Thus, a novel token merging method is required to improve the computational efficiency of token processing within well-trained video generation models, while preserving high video quality.
@@
C3) High Memory Requirements. Fig. 2 shows the memory requirements of model components in Open-Sora [83], i.e., VAE [15], T5 [48], and STDiT [83], where their cumulative memory demand, i.e., 23 GB, surpasses the memory capacity of many mobile devices. For instance, the iPhone 15 Pro [1], with 8 GB of memory, restricts the available memory for a single application to 3.3 GB for system stability. Furthermore, the individual memory requirements of T5 and STDiT exceed 3.3 GB, creating challenges in loading them into memory. In addition, some memory must be reserved for model execution (inference), exacerbating memory shortages on mobile devices. Thus, limited device memory is another challenge that should be addressed to enable on-device video generation.
@@
Figure 2: The size of Open-Sora models: T5 [48] (18.00 GB), STDiT [83] (4.50 GB), and VAE [15] (0.82 GB), which exceeds the available memory capacity of the iPhone 15 Pro [1] (3.3 GB).
3 Overview: On-device Sora
Fig. 1 shows an overview of On-Device Sora, which takes Open-Sora [83] outlined in Sec. 2.1 as its backbone. On-device Sora enables efficient diffusion-based text-to-video generation on mobile devices by addressing three key challenges (Sec. 2.2) through three proposed methods: 1) Linear Proportional Leap (LPL) (Sec. 4), 2) Temporal Dimension Token Merging (TDTM) (Sec. 5), and 3) Concurrent Inference with Dynamic Loading (CI-DL) (Sec. A.1). These methods can also be applied to other diffusion-based text-to-video generation models, e.g., Pyramidal Flow [29].
4 Linear Proportional Leap